The New England cheese company produces two cheeses. Company produces two cheese 

spread by blending mild cheddar cheese with extra sharp cheddar cheese. The cheese spreads 

are packaged in 12-ounce container which are then sold to distributors throughout the 

Northeast. The regular blend contains 80% mild cheddar and 20% extra sharp and the Zysty 

blend contains 60% mild cheddar and 40% extra sharp. This year a local dairy cooperative 

offered to provide up to 8100 pounds of mild cheddar cheese $1.20 per pound and up to 3000 

pounds of extra sharp cheddar cheese for $ 1.40 per pound. The cost to blend and package the 

cheese spreads excluding the cost of cheese is $0.20 per container. If each container of regular 

is sold for $1.95 and each container of Zesty is sold for dollar 2.20 how many containers off 

regular and justice should New England cheese produce. Formulate the above problem into a 

linear programming problem and solve it using Excel Solver. (5)


1
Expert's answer
2022-01-10T12:46:01-0500

n your constraints you forgot that each container is .75 lb


Regular has

.80*.75 = 0.6 lb mild

.20*.75 = 0.15 lb sharp


Zesty has

.60*.75 = 0.45 lb mild

.40*.75 = 0.3 lb sharp


Cost per container is:


R: 0.6*1.2 + 0.15*1.4 + 0.2 = 1.13

Z: 0.45*1.2 + 0.3*1.4 + 0.2 = 1.16


So, total profit is


P = R*(1.95 - 1.13) + Z*(2.20 - 1.16)


Constraints:

0.6*R + 0.45*Z <= 8100

0.15*R + 0.3*Z <= 3000


Regular has

.80*.75 = 0.6 lb mild

.20*.75 = 0.15 lb sharp


Zesty has

.60*.75 = 0.45 lb mild

.40*.75 = 0.3 lb sharp


Cost per container is:


R: 0.6*1.2 + 0.15*1.4 + 0.2 = 1.13

Z: 0.45*1.2 + 0.3*1.4 + 0.2 = 1.16


So, total profit is


P = R*(1.95 - 1.13) + Z*(2.20 - 1.16)


Constraints:

0.6*R + 0.45*Z <= 8100

0.15*R + 0.3*Z <= 3000